OpenStudy (anonymous):

@paki Check this? :) Sphere Volume Formula V = 4/3 x 3.14 x r^3 A sphere has a radius of 6 cm. What is the volume of the sphere? A. 216 cm^3 B. 288 cm^3 C. 216π cm^3 D. 288π cm^3 R = 6^3 = 216 How do I do this? Can you help me without giving me the answer? o_o

OpenStudy (paki):

V = 4/3 x 3.14 x 216

OpenStudy (paki):

solve now... and also dont get confused in the value of 4/3....

OpenStudy (paki):

or use V = 1.3333 x 3.14 x 216

OpenStudy (anonymous):

Do I do (3/4) x 3.14 x 216?

OpenStudy (paki):

rite

OpenStudy (paki):

good... now you got the point... :)

OpenStudy (anonymous):

I put it into my calculator and got... 881.712 Now what do I do? :)

OpenStudy (paki):

no.... i got 904.09

OpenStudy (anonymous):

Hm... I put it in like this.. 1.3 x 3.14 x 216

OpenStudy (paki):

no... its 1.333 not 1.3

OpenStudy (anonymous):

Oh! OK let me try that again. :)

OpenStudy (paki):

yeah... try again...

OpenStudy (anonymous):

OK, I got what you got now. :) I see what I did wrong. OK so NOW what do we do next? :)

OpenStudy (anonymous):

its 904.09 now multiply as well as divide it bye pie and u'll get 288π cm^3

OpenStudy (anonymous):

*by

OpenStudy (paki):

cool :)

OpenStudy (anonymous):

OK, Now I'm confused again. You want me to multiply 904.09 by 3.14, which equals 2838.8426 and divide it by what?

OpenStudy (paki):

909.09 is the answer....

OpenStudy (paki):

no need to multiply and divide it by pie value... we have used pie value already in the formula...

OpenStudy (anonymous):

no divide it by 3.14 and just put the pie sign with the ans like this 288π

OpenStudy (anonymous):

@paki the ans choices require that

OpenStudy (paki):

no answer choices are wrong...

OpenStudy (anonymous):

The exact question is... A sphere has a radius of 6 cm. What is the volume of the sphere? A. 216 cm^3 B. 288 cm^3 C. 216π cm^3 D. 288π cm^3

OpenStudy (anonymous):

So...D?

OpenStudy (paki):

have a look here.... http://calculator.tutorvista.com/sphere-calculator.html

OpenStudy (anonymous):

ok....

OpenStudy (paki):

yeah...

OpenStudy (paki):

all the values which you need is over there... @CrazyCountryGirl

OpenStudy (anonymous):

kk, hold on. brb! :)

OpenStudy (paki):

ok be quick... i am going to log off... as here it's 12:36am...

OpenStudy (anonymous):

OK, I'm back. :) It's only 3'37 pm here. :)

OpenStudy (paki):

hmmm here now 12:38AM... at pakistan....

OpenStudy (anonymous):

The thing says that the volume 904.779. I just divided it by 3.14 and got 288.

OpenStudy (anonymous):

I'm in the us. :)

OpenStudy (paki):

hmm... but dividing it by 3.14 makes no sense.... :/

OpenStudy (paki):

let me tag @ganeshie8 please help here...

OpenStudy (paki):

@ganeshie8 please check this question....?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (paki):

@CrazyCountryGirl i am leaving ok... best of luck for your question....

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

Thanks! @ganeshie8 @Harsha19111999 @geerky42 @mathmate @mathstudent55 @myininaya @Luigi0210 @tHe_FiZiCx99 @Preetha @nincompoop

ganeshie8 (ganeshie8):

use the formula : Volume of sphere = \(\large \dfrac{4}{3}\pi r^3\)

ganeshie8 (ganeshie8):

simply plugin \(\large r = 6\)

OpenStudy (anonymous):

OK. :) I'm pretty sure R=6^3 though. :)

ganeshie8 (ganeshie8):

Volume of sphere = \(\large \dfrac{4}{3}\pi r^3 = \large \dfrac{4}{3}\pi (6)^3 = \large \dfrac{4}{3}\pi (216) = 4 \pi (72) = 288 \pi\)

OpenStudy (anonymous):

I see! Thanks! :)

ganeshie8 (ganeshie8):

yw :)